Please try to give the correct answer with the detailed description of how you solved. Thanks.Reference Electrodes The Eº value for the following reaction is 0.446 V relative to the standard hydrogen electrode (SHE): Ag
Answer #1

Can Given cell Reaction Ag,croq (s1 & 26 2 Agisl + crop lag E = stand and Reduction potential for Half cell = Esce = o. 242 v0 [cro2 = 2.60*108M Ecce = Focu - oooss log , T = 298k anges ² x crop 2 - 1 OAB. Creast 1 x = 2.60X - 0. zoavelt - oss log (Гcto42-1 = 4.0 x 10 M TAngо е